### v1

@@ Abstract
Previous fake speech datasets were constructed from a defender’s perspective to develop countermeasure (CM) systems without considering diverse motivations of attackers. To better align with real-life scenarios, we created LlamaPartialSpoof, a 130-hour dataset contains both fully and partially fake speech, using a large language model (LLM) and voice cloning technologies to evaluate the robustness of CMs. By examining information valuable to both attackers and defenders, we identify several key vulnerabilities in current CM systems, which can be exploited to enhance attack success rates, including biases toward certain text-to-speech models or concatenation methods. Our experimental results indicate that current fake speech detection system struggle to generalize to unseen scenarios, achieving a best performance of 24.44% equal error rate.
@@ II-A Motivation
To align with potential real-life threats, we created the LlamaPartialSpoof dataset which includes both fully and partially fake speech utterances. Unlike the previous databases [15, 20, 16], ours was generated using a pipeline designed from attackers’ perspective. It prioritizes on content, intention, and quality of the synthetic speech. Since several studies have shown training detection models on simulated data can improve performances[20, 21, 22], it is not necessary for the training data to be semantically correct. Hence, we focus on providing a high-quality, realistic, and diverse evaluation set. The generation pipeline presented in this paper can be replicated to create training data under similar conditions.
@@ II-B Generation Pipeline
TTS001 LJ JET 5,580 5,414 2.41 4.69
TTS002 YourTTS 5,580 5,414 5.45 6.64
TTS003 XTTS V2 5,580 5,411 2.40 4.21
TTS004 GPT Sovits 5,580 5,414 4.59 5.33
TTS005 CosyVoice 5,579 5,288 4.57 4.37
TTS006 ElevenLab 5,580 5,414 1.82 3.20
@@
Ours LlamaPartialSpoof LibriTTS [23] EN/Yes Crossfade 6 10,573 33,479 32,355
@@ II-C Text-to-speech Models
LJ JET: This single female speaker TTS model [29] was trained on the 24-hour LJ Speech dataset [^4] https://keithito.com/LJ-Speech-Dataset/. It [^5] https://github.com/espnet/espnet/tree/master/egs2/ljspeech/tts1 is an end-to-end model that jointly training the FastSpeech 2 feature generator [30] and the HiFi-GAN neural vocoder [31]. We used it as a non-voice-cloning reference.
@@ II-D Dataset Analysis
We created 5,580 spoofed sentences out of 5,736 originals after checking outputs of the Llama model. Among those, 5,414 sentences were used to create partially fake utterances, excluding 166 sentences only had delete operations. Table I shows the number of fully and partially fake utterances created with each TTS model [^11] TTS005 has fewer samples due to failed alignment in some cases. We calculated word error rate (WER) on 5,000 sentences using Whisper large model [^12] https://huggingface.co/openai/whisper-large. Overall, partial subsets had higher WER than their full counterparts with ElevenLab got the best results. Figure 2 shows the distributions of fake sentences across different percentage of fake words or different number of fake segments. Comparing with the AV-Deepfake1M dataset, our partial subset has more utterances with more than two fake segments. The shortest fake segment is 0.03 s, while the longest is 24.72 s, with a mean of 0.96 s and a standard deviation of 0.87 s.
@@
A. ps-train 1.48 40.40 39.37 5.14 24.44
B. had-train 49.51 0.02 49.81 42.24 57.36
C. 1m-train 51.08 56.29 0.02 51.41 51.70
D. asv5-train 30.73 33.22 43.99 4.50 33.43
A+B+C 1.19 0.06 0.04 14.07 32.14
@@
A. ps-train 13.70 47.41 20.59 5.81 46.17
B. had-train 43.97 0.12 28.61 35.56 59.52
C. 1m-train 49.84 40.76 0.18 60.60 50.42
D. asv5-train 34.20 37.74 34.99 2.99 37.92
A+B+C 13.59 0.08 0.19 26.46 47.63
@@ III Experiments
Train Set LJ JET YourTTS XTTS V2 GPT-SoVITS CosyVoice ElevenLab All
@@
A. ps-train 47.32 17.22 20.05 10.66 27.04 14.76 18.36 13.71 30.36 16.56 46.04 18.08 32.36 15.23
B. had-train 45.00 36.74 66.76 51.00 70.44 53.16 68.02 55.08 66.42 48.04 69.34 48.52 65.14 48.45
C. 1m-train 55.27 48.97 54.07 48.13 55.54 49.63 54.74 50.52 54.15 48.64 51.18 47.33 54.18 48.95
D. asv5-train 45.82 43.91 0.89 12.47 24.82 26.66 8.67 21.40 40.23 36.98 48.77 44.71 33.81 33.61
A+B+C 48.65 28.40 18.27 17.36 32.93 21.04 40.19 24.85 47.86 27.49 38.80 26.36 38.97 24.41
@@ III-D Results and Analysis
A. ps-train 22.39 13.10 10.85 12.99 18.75
B. had-train 44.85 49.06 48.88 49.04 52.15
C. 1m-train 50.18 47.64 48.49 49.92 52.07
D. asv5-train 40.42 35.01 27.94 27.82 30.06
A+B+C 28.72 23.97 21.70 23.23 26.49
@@
A. ps-train 32.10 15.11 15.20 13.93
B. had-train 64.63 48.26 32.31 38.33
C. 1m-train 53.79 49.11 55.25 54.07
D. asv5-train 33.65 33.17 33.80 29.09
A+B+C 38.73 24.34 21.25 24.31
@@ III-D1 Cross-dataset fake speech detection
Table III(b) presents the utterance-based and 20-ms segment-based EER when training and evaluating across datasets. The model A+B+C was trained using the combination of three partially fake training sets. As expected, both types of EERs are every low when the train and evaluation set come from the same source but they are very high otherwise. Model A trained with ps-train had the best utterance-based score while model D trained with asv5-train had the best segment-based when tested with the LlamaPartialSpoof dataset. We believe The diverse range of attacks in the training sets helped these models generalize better to unseen ones. The degradation performances when training and testing across datasets highlight the need to develop better strategies for domain adaption and generalization.
@@ III-D2 EERs of fully and partially fake subsets
Table IV shows the EER evaluating on the partial and full subsets generated by different TTS models. Generally speaking, model A is better at detecting partially fake than fully fake while it is the opposite for model D due to the nature of their training data. LJ JET had the highest lowest-EER for the fully fake scenario while it is ElevenLab for the partially fake. Interestingly, partially fake samples of LJ JET got the 2nd highest EER which suggests the CMs did not take speaker consistency into account. Model D, trained with ASVspoof5 data, can detect YourTTS fully fakes samples quite well as its data also includes samples created with YourTTS [7]. Since model A, trained on the PartialSpoof dataset, had the best performance on detecting partial subsets, it reiterates that semantically ignorant data can be used to train partially fake speech detection model.
@@ III-D4 Effects of concatenating methods
Table V(b) shows the utterance-based EER calculated with fully fake or partially fake samples created with different methods. Model A trained with PartialSpoof, which used cross-correlation and overlap-add, is better at detecting samples created with the same method. Interestingly, model B and C are better at detecting samples created with cut/paste and crossfade respectively, but their EERs are still very high. Model D favors samples created with overlap-add even though its training data does not contain partially fake samples. This highlights concatenation methods as an aspect of interest for both attackers and defenders.
@@ IV Conclusion
In this paper, we show how attackers can utilize LLMs and voice cloning technologies to create mass disinformation media. Researchers can also use this framework to conduct experiments on deepfake creation and detection across transcript and audio [2]. Our 130-hour LlamaPartialSpoof dataset is an initiative to create a large-scale out-of-domain evaluation set for CM research and development. Several aspects of creating partially fake speech were investigated, but the overall methodology was still very simple. For the future work, researchers can focus on improving performances of CM systems on these scenarios or more advanced ones [35] such as in noisy [12] or reverberant environments.

### v2

@@ Abstract
Previous fake speech datasets were constructed from a defender’s perspective to develop countermeasure (CM) systems without considering diverse motivations of attackers. To better align with real-life scenarios, we created LlamaPartialSpoof, a 130-hour dataset that contains both fully and partially fake speech, using a large language model (LLM) and voice cloning technologies to evaluate the robustness of CMs. By examining valuable information for both attackers and defenders, we identify several key vulnerabilities in current CM systems, which can be exploited to enhance attack success rates, including biases toward certain text-to-speech models or concatenation methods. Our experimental results indicate that the current fake speech detection system struggle to generalize to unseen scenarios, achieving a best performance of 24.49% equal error rate.
@@ II-A Motivation
To align with potential real-life threats, we created the LlamaPartialSpoof dataset, which includes both fully and partially fake speech utterances. Unlike the previous databases [15, 20, 16], ours was generated using a pipeline designed from attackers’ perspective. It prioritizes content, intention, and quality of synthetic speech. Since several studies have shown that training detection models on simulated data can improve performance[20, 21, 22], it is not necessary for the training data to be semantically correct. Hence, we focus on providing a high-quality, realistic, and diverse evaluation set. The generation pipeline presented in this paper can be replicated to create training data under similar conditions.
@@ II-B Generation Pipeline
TTS001 LJ JETS 5,577 5,387 2.28 3.59
TTS002 YourTTS 5,577 5,387 5.24 5.52
TTS003 XTTS V2 5,577 5,384 2.29 3.42
TTS004 GPT Sovits 5,577 5,387 4.13 4.23
TTS005 CosyVoice 5,576 5,286 3.23 3.50
TTS006 ElevenLab 5,577 5,387 1.27 2.39
@@
Ours LlamaPartialSpoof LibriTTS [23] EN/Yes Crossfade 6 10,573 33,461 32,194
@@ II-C Text-to-speech Models
LJ JETS: This single female speaker TTS model [29] was trained on the 24-hour LJ Speech dataset [^4] https://keithito.com/LJ-Speech-Dataset/. It [^5] https://github.com/espnet/espnet/tree/master/egs2/ljspeech/tts1 is an end-to-end model that jointly training the FastSpeech 2 feature generator [30] and the HiFi-GAN neural vocoder [31]. We used it as a non-voice-cloning reference.
@@ II-D Dataset Analysis
We created 5,577 spoofed sentences out of 5,736 originals after checking outputs of the Llama model. Among those, 5,387 sentences were used to create partially fake utterances, excluding 190 sentences only had delete operations. Table I shows the number of fully and partially fake utterances created with each TTS model [^11] TTS005 has fewer samples due to failed alignment in some cases. We calculated word error rate (WER) on 5,000 sentences using Whisper large model [^12] https://huggingface.co/openai/whisper-large. Overall, partial subsets had higher WER than their full counterparts with ElevenLab got the best results. Figure 2 shows the distributions of fake sentences across different percentage of fake words or different number of fake segments. Comparing with the AV-Deepfake1M dataset, our partial subset has more utterances with more than two fake segments. The shortest fake segment is 0.03 s, while the longest is 24.72 s, with a mean of 0.95 s and a standard deviation of 0.85 s.
@@
A. ps-train 1.48 40.40 39.37 5.14 24.49
B. had-train 49.51 0.02 49.81 42.24 57.18
C. 1m-train 51.08 56.29 0.02 51.41 51.63
D. asv5-train 30.73 33.22 43.99 4.50 34.21
A+B+C 1.19 0.06 0.04 14.07 31.87
@@
A. ps-train 13.70 47.41 20.59 5.81 47.49
B. had-train 43.97 0.12 28.61 35.56 59.36
C. 1m-train 49.84 40.76 0.18 60.60 51.16
D. asv5-train 34.20 37.74 34.99 2.99 38.50
A+B+C 13.59 0.08 0.19 26.46 50.16
@@ III Experiments
Train Set LJ JETS YourTTS XTTS V2 GPT-SoVITS CosyVoice ElevenLab All
@@
A. ps-train 47.31 17.14 20.25 10.72 27.11 14.90 18.47 13.50 30.62 16.92 46.13 18.37 32.35 15.33
B. had-train 44.93 36.02 66.55 49.65 70.58 52.41 68.21 53.56 66.37 47.50 69.35 47.87 65.21 47.54
C. 1m-train 55.23 48.99 54.12 48.26 55.68 49.58 54.79 50.09 54.05 48.22 51.21 47.36 54.17 48.90
D. asv5-train 44.98 44.10 0.63 9.03 24.25 26.33 6.60 22.63 40.17 37.64 48.41 45.29 34.14 34.52
A+B+C 50.11 27.84 18.39 17.24 32.61 20.83 40.07 24.80 47.88 27.35 38.75 26.57 39.00 24.55
@@ III-D Results and Analysis
A. ps-train 22.80 13.51 11.12 12.76 19.42
B. had-train 44.37 47.59 47.56 47.84 52.16
C. 1m-train 50.43 47.41 48.07 50.10 51.33
D. asv5-train 41.03 36.05 27.14 27.15 31.66
A+B+C 28.78 24.21 21.54 22.68 26.68
@@
A. ps-train 32.07 15.21 15.37 13.89
B. had-train 64.78 47.24 32.72 37.89
C. 1m-train 53.66 49.04 54.90 54.06
D. asv5-train 34.17 34.26 35.11 28.56
A+B+C 38.70 24.54 21.83 24.21
@@ III-D1 Cross-dataset fake speech detection
Table III(b) presents the utterance-based and 20-ms segment-based EER when training and evaluating across datasets. The model A+B+C was trained using the combination of three partially fake training sets. As expected, both types of EERs are every low when the train and evaluation set come from the same source, but they are very high otherwise. Model A trained with ps-train had the best utterance-based score while model D trained with asv5-train had the best segment-based when tested with the LlamaPartialSpoof dataset. We believe The diverse range of attacks in the training sets helped these models generalize better to unseen ones. The degradation performances when training and testing across datasets highlight the need to develop better strategies for domain adaption and generalization.
@@ III-D2 EERs of fully and partially fake subsets
Table IV shows the EER evaluating on the partial and full subsets generated by different TTS models. Generally speaking, model A is better at detecting partially fake than fully fake while it is the opposite for model D due to the nature of their training data. LJ JETS had the highest lowest-EER for the fully fake scenario while it is ElevenLab for the partially fake. Interestingly, partially fake samples of LJ JETS got the 2nd highest EER which suggests the CMs did not take speaker consistency into account. Model D, trained with ASVspoof5 data, can detect YourTTS fully fakes samples quite well as its data also includes samples created with YourTTS [7]. Since model A, trained on the PartialSpoof dataset, had the best performance on detecting partial subsets, it reiterates that semantically ignorant data can be used to train partially fake speech detection model.
@@ III-D4 Effects of concatenating methods
Table V(b) shows the EER based on utterance calculated with fully or partially fake samples created with different methods. Model A trained with PartialSpoof, which used cross-correlation and overlap-add, is better at detecting samples created with the same method. Interestingly, model B and C are better at detecting samples created with cut/paste and crossfade respectively, but their EERs are still very high. Model D favors samples created with overlap-add even though its training data does not contain partially fake samples. This highlights concatenation methods as an aspect of interest for both attackers and defenders.
@@ IV Conclusion
In this paper, we show how attackers can utilize LLMs and voice cloning technologies to create mass disinformation media. Researchers can also use this framework to conduct experiments on deepfake creation and detection across transcripts and audio [2]. Our 130-hour LlamaPartialSpoof dataset is an initiative to create a large-scale out-of-domain evaluation set for CM research and development. Several aspects of partially fake speech creation were investigated, but the overall methodology was still very simple. For the future work, researchers can focus on improving performances of CM systems on these scenarios or more advanced ones [35] such as in noisy [12] or reverberant environments [36].